OpenStudy (nuke_ur_friend):

What is the product of eight over fifteen times ten over eleven in simplest form? sixteen over thirty-three nine over thirteen eighty-eight over one hundred fifty sixteen over one hundred sixty-five @turtleluver

OpenStudy (anonymous):

16/33

OpenStudy (texaschic101):

8/15 x 10/11 = 80/165 = 16/33.....I agree with turtleluver :)